OpenStudy (anonymous):

e^0=1 e^1=0 why is that so?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

cuz if u ln both side ur ans would be a zero

OpenStudy (anonymous):

well its commutative property thats when you swap the nmbers examle below: a+b=c a+c=b

OpenStudy (anonymous):

e^1 = e

OpenStudy (anonymous):

also the property of zero

OpenStudy (amistre64):

yeah, um... e^1 = e :)

OpenStudy (amistre64):

e^1 = 2.71828182845905...

OpenStudy (anonymous):

and every number ^ 0 is 1.

OpenStudy (anonymous):

2.718... = e